OpenStudy (anonymous):

What is the value of x in the equation 2 ? –24 –12 12 24

OpenStudy (anonymous):

equation

jimthompson5910 (jim_thompson5910):

Is something missing (like a coefficient outside the matrix)?

OpenStudy (anonymous):

a 2 is outside the matrix..sorry the picture didn't include it

jimthompson5910 (jim_thompson5910):

alright thank you

jimthompson5910 (jim_thompson5910):

Double the upper right value in the first matrix: 2*9 = 18 Then subtract off -6 (the corresponding upper right value in the second matrix) 18 - (-6) = 18+6 =24 So x = 24

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

you're welcome